OpenStudy (pokemonfan40):

Which equation can be used to find the answer? Camillia needs $65 for a pair of hiking shoes. she has already $19 saved. How much more does she need to save? A n + 19 = 65 n = 46 $46 B n -:- 2 = 19 n = 38 $38 C n - 19 = 65 n = 84 $84 D n -:- 2 = 65 n = 130 $130

OpenStudy (mynameisnemo):

a :)

OpenStudy (pokemonfan40):

thanks!

OpenStudy (mynameisnemo):

would you like me to explain why?

OpenStudy (pokemonfan40):

yeah

OpenStudy (mynameisnemo):

No? okie

OpenStudy (mynameisnemo):

oh okay

OpenStudy (mynameisnemo):

the reason why the answer is A is because in order to find out how much money she needs to save, you need to get the variable (n) alone. N stands for the money needed. you need to isolate n by subtracting 19 from both sides, the equation being: n + 19 = 65 - 19 - 19 n = 46 at this point, you're left with: money needed to buy shoes = $46

OpenStudy (pokemonfan40):

thanks man here is a medal!

OpenStudy (mynameisnemo):

haha thank you, soon I will become the best trainer ever and my goal is to have caught all of thee pokemon! Your medal really helped